# Giovanni Compiani Ilya Morozov Stephan Seiler

### v3

@@ Appendix F Changes in Text and Images Over Time
Book Plain Logit Mixed Logit with Attributes Mixed Logit with Texts
Don’t Believe 3.7% 6.7% 8.9%
Art of Letting Go 3.0% 4.3% 5.6%
The Ritual 3.1% 1.6% 2.5%
The Inmate 5.0% 4.4% 4.3%
Ashes & Star 2.7% 1.3% 2.0%
Court of Ravens 3.3% 3.5% 2.5%
Serpent & Wings 3.4% 2.3% 2.3%
Please Tell Me 6.9% 6.1% 5.6%
The Housemaid 5.4% 5.2% 4.0%
Table A2: Merger Simulation Results (Dopamine Detox). The table shows predicted relative price increases resulting from a simulated merger between Dopamine Detox and each other book respectively. For each simulated merger between Dopamine Detox and another book (first column), the table reports the expected average price increase for the two merging books across three estimated demand models.
@@
[…]
Category: Tablets
Mixed Logit with Attributes 4293.2 -54.9
Mixed Logit with Unstructured Data 4236.6 -111.5
Category: Monitors
Mixed Logit with Attributes 1376.4 -12.3
Mixed Logit with Unstructured Data 1366.4 -22.4
Category: Memory Cards
Mixed Logit with Attributes 2709.3 -72.6
Mixed Logit with Unstructured Data 2685.1 -96.8
Category: Headphones
Mixed Logit with Attributes 9882.0 -4.5
Mixed Logit with Unstructured Data 9875.8 -10.7
Table A4: AICAIC Comparison of Our Approach and Mixed Logit with Attributes
Fire 7 Fire HD Fire HD Fire 7 iPad Fire HD Dragon iPad
8 10 Kids 10.2 8 Kids Touch 9.7
Fire 7 0.000 0.330 0.381 0.301 0.312 0.283 0.279 0.286
Fire HD 8 0.225 0.000 0.229 0.175 0.178 0.165 0.162 0.166
Fire HD 10 0.386 0.340 0.000 0.314 0.320 0.297 0.291 0.295
Fire 7 Kids 0.123 0.105 0.128 0.000 0.100 0.091 0.089 0.091
iPad 10.2 0.153 0.129 0.148 0.122 0.000 0.113 0.112 0.114
Fire HD 8 Kids 0.043 0.038 0.049 0.035 0.035 0.000 0.032 0.033
Dragon Touch 0.022 0.018 0.021 0.017 0.017 0.016 0.000 0.016
iPad 9.7 0.048 0.041 0.044 0.037 0.037 0.035 0.034 0.000
Fire 7 Fire HD Fire HD Fire 7 iPad Fire HD Dragon iPad
8 10 Kids 10.2 8 Kids Touch 9.7
Fire 7 0.000 0.590 0.162 0.362 0.143 0.033 0.024 0.080
Fire HD 8 0.753 0.000 0.429 0.336 0.200 0.193 0.108 0.116
Fire HD 10 0.058 0.144 0.000 0.079 0.142 0.297 0.022 0.077
Fire 7 Kids 0.127 0.131 0.080 0.000 0.012 0.440 0.293 0.007
iPad 10.2 0.049 0.079 0.141 0.010 0.000 0.015 0.310 0.715
Fire HD 8 Kids 0.003 0.034 0.151 0.200 0.004 0.000 0.188 0.001
Dragon Touch 0.000 0.001 0.000 0.011 0.013 0.019 0.000 0.003
iPad 9.7 0.011 0.020 0.036 0.002 0.485 0.003 0.055 0.000

### v4

@@ Appendix G Changes in Text and Images Over Time
Fire 7 Fire HD Fire HD Fire 7 iPad Fire HD Dragon iPad
8 10 Kids 10.2 8 Kids Touch 9.7
Fire 7 0.000 0.330 0.381 0.301 0.312 0.283 0.279 0.286
Fire HD 8 0.225 0.000 0.229 0.175 0.178 0.165 0.162 0.166
Fire HD 10 0.386 0.340 0.000 0.314 0.320 0.297 0.291 0.295
Fire 7 Kids 0.123 0.105 0.128 0.000 0.100 0.091 0.089 0.091
iPad 10.2 0.153 0.129 0.148 0.122 0.000 0.113 0.112 0.114
Fire HD 8 Kids 0.043 0.038 0.049 0.035 0.035 0.000 0.032 0.033
Dragon Touch 0.022 0.018 0.021 0.017 0.017 0.016 0.000 0.016
iPad 9.7 0.048 0.041 0.044 0.037 0.037 0.035 0.034 0.000
Fire 7 Fire HD Fire HD Fire 7 iPad Fire HD Dragon iPad
8 10 Kids 10.2 8 Kids Touch 9.7
Fire 7 0.000 0.939 0.451 0.144 0.328 0.019 0.347 0.073
Fire HD 8 0.592 0.000 0.027 0.023 0.068 0.002 0.267 0.008
Fire HD 10 0.247 0.023 0.000 0.167 0.213 0.328 0.299 0.332
Fire 7 Kids 0.035 0.006 0.090 0.000 0.055 0.624 0.000 0.016
iPad 10.2 0.100 0.025 0.147 0.072 0.000 0.018 0.058 0.558
Fire HD 8 Kids 0.004 0.000 0.174 0.582 0.009 0.000 0.000 0.011
Dragon Touch 0.003 0.004 0.006 0.000 0.001 0.000 0.000 0.002
iPad 9.7 0.018 0.002 0.105 0.012 0.326 0.009 0.029 0.000
Table A3: Estimated diversion ratios for tablets (mixed logit with observed attributes). Each cell reports the probability of choosing the row product $j$ when the first choice, column product $k$, is removed from the choice set.
Fire 7 Fire HD Fire HD Fire 7 iPad Fire HD Dragon iPad
8 10 Kids 10.2 8 Kids Touch 9.7
Fire 7 0.000 0.590 0.162 0.362 0.143 0.033 0.024 0.080
Fire HD 8 0.753 0.000 0.429 0.336 0.200 0.193 0.108 0.116
Fire HD 10 0.058 0.144 0.000 0.079 0.142 0.297 0.022 0.077
Fire 7 Kids 0.127 0.131 0.080 0.000 0.012 0.440 0.293 0.007
iPad 10.2 0.049 0.079 0.141 0.010 0.000 0.015 0.310 0.715
Fire HD 8 Kids 0.003 0.034 0.151 0.200 0.004 0.000 0.188 0.001
Dragon Touch 0.000 0.001 0.000 0.011 0.013 0.019 0.000 0.003
iPad 9.7 0.011 0.020 0.036 0.002 0.485 0.003 0.055 0.000